OpenStudy (anonymous):

What is the slope of the line passing through the points (–3, 4) and (2, –1)? A. 07_04_LQ_3c.gif B. –1 C. 07_04_LQ_3d.gif D. 1

OpenStudy (anonymous):

what are the answer choices for A and C?

OpenStudy (anonymous):

go with " 07_04_LQ_3d.gif " because C is usually correct

OpenStudy (igreen):

\(m = \dfrac{-1-4}{2+3}\) Simplify

OpenStudy (anonymous):

^^^ iGreen is correct. follow those steps

OpenStudy (anonymous):

A. 3/5

OpenStudy (chosenmatt):

yes what @iGreen said

OpenStudy (anonymous):

C -5/3

OpenStudy (chosenmatt):

:( why am i always the last person to say something xD

OpenStudy (anonymous):

from -3 to 2 is right five units from 4 to -1 is down five right five, down five, same as right one down one, slope is \(-1\)

OpenStudy (igreen):

Oh wait, you were just telling us what the options were :P

OpenStudy (anonymous):

yes

OpenStudy (igreen):

@satellite73 You're not supposed to give answers..

OpenStudy (anonymous):

or you can do it the math teacher way \[\frac{4-(-1)}{3-(-2)}\] yeah but at at least i gave an explanation, not a slave to a formula

OpenStudy (igreen):

You're still not supposed to give answers, explanation or no.

OpenStudy (chosenmatt):

um...i think im gonna go....bye....lol

OpenStudy (igreen):

Lol

OpenStudy (anonymous):

trying to make some sense out of what it means any donkey can compute \[\frac{y_2-y_1}{x_2-x_1}\] but it takes a human being to understand that it means the vertical direction up or down for a single increase in x